APPRENTICE (and possible SIT-INS)

Name (32)  @uchicago.edu  Courses  other topics
William Cole Abram  abramwc  5, 2c, 4b, 3a, 3c, 3b, 2a  combinatorics, number thy
Dylan Allegretti  allegretti  5, 4a, 3a, 1  logic, set theory, model thy
Adam Anderson  adamanderson  5, 4a, 2b, 2a, 1, 3  analysis, math. phys, chaos
Elan Bechor  ebechor  5, 1, 3a, 2a, 2b  information thy, axiomatic set thy
Michael Calderbank  Mcalderbank  5, 3a, 2a, 2b, 4a  number thy, polytopes, info thy
Daniel Cervone (??)  dcervone  5, 1, 2b, 4a, 2a
Beau Dabbs  bdabbs13  5, 2, 4, 1, 3
Catherine Easton  catherine  5, 1, 2b, 4a
Mufti Nomaan Farooqui  mnomy88  5, 1, 2c, 2a, 3b  matrices
Hannah Gordon  hannahg  5. 2c. 1  music
Austin Gross  Agross  5, 1, 2a, 3a, 4a, 2b  topology?
Eric Daniel Jaffe  ericjaffe  5, 1, 2b, 4a, 3a  number thy, integrals
Ethan Jerzak  jerzak  5, 2a, 1, 2c, 4b, 3a
Michael (Ryan) Julian  mrjulian  5, 2b, 4a, 4b, 3a  diff equ, dynamics, topology
Katayun J. Kamdin  kkamdin  5, 2a, 2b, 4b, 3b
Satory Kudose  skudose  5, 1, 3abc, 4ab, 2abc  number theory
William Leeb  weleeb  5, 2, 1, 3  set theory
Chenchuan Li  chenchuan  5, 3a, 2a, 3b, 4a, 4b  probability, analysis
Cong Han Lim  conghan  5, 1, 2
Ellen McCullagh  ellenm  1, 3a, 4a  diff. geom, knot theory, topology
Scott Messick  smessick  5, 2, 4, 1, 3  (integrals)
Currence Monson  currence  5, 1, 3a, 3b
Matt Olson  molson2  5, 2a, 2b  measure theory and probability
Bo Peng  bopeng  1, 5, 2c, 4a, 2a, 3a  random walks
Andrew Poppick  apoppick  5, 2b, 3a
James Salvatore  jsal13  1, 2a, 2b, 3a  algebra, topology of any kind
Honi Sanders  honiraphael  5, 1, 2a, 3, 4a
Brad Sugarman  bvs  5, 3a, 3b, 4a, 4b
Michael Travis  mtravis  5, 2c, 2a, 4b, 2b, 1
Brooke Ullery  bullery  5, 1, 2a, 3a, 3b
Gideon Tan Guang Yuan  serapuff  2, 1, 3a, 4a, 5  Fin math; stoch proc
Tengren Zhang  tengren85  2a, 1, 5, 3a, 4a  group theory

SESAME

Name (22)  @uchicago.edu  preferences  other topics
Stephan Balady  sbalady  4a, 1, 3a, 4b, 2b  Haar measure; analysis or topology
Sarah Bennett  sbennett  2c, 4b, 3b, 3a, 3c
David Berger  iloveyou  2a, 3a, 1, 4a, 2c
Chad Casarotto  chadmbol  1, 2a  discrete math or probability theory
Chris Logan Chariker  chariker  3c, 2a, 4a, 4b, 2b, 1, 3b  Galois theory
Le Chen  lelechen  1, 2a, 2b, 2c, 4a  real analysis, probability theory
David Coley  dcoley  2b, 3c, 5(?)  number theory, algebraic geometry
Andrew Eisenberg  aeisen  3a, 4a, 4b, 1, 3b, 2a  algebra
Arpit Gupta  arpit  2a, 1, 2b
Adam Kaye  adamkaye  3abc, 2a, 4ab  maybe algebraic number theory
R. Andrew Kreek  rakreek  1, 2a, 2b, 3b  geometry in general
Joshua Blaise Leners  leners  1, 2a, 3c, 4b, 2c, 4a, 2b, 3a, 3b GRE math, complexity, game thy
Carmel Levy  carmel1  1, 3b, 3a, 4a, 2a, 2b, 2c, 4b, 5  complexity theory
Yana Peysakhovich  yana  1, 2a, 2b, 3b, 3a, 4a  1 (that sounds good)
Robert Prag  rprag  2b, 1, 2a, 4b, 4a, 3b, 3c  rings, modules, ⊗, alg. number thy
Edward Ruggeri  ruggeri  3a, 2a, 4b, 3b, 3c  topology, graphs
Maxwell Shron  mshron  3a, 3b, 1, 2a, 2b  hyperbolic geom, low dim top
Andries Smith  andries  1, 4, 3, 2, 5
Michael Thvedt  cynicalk  4a, 4b, 3a, 3b, 2a, 2b  music, TQFT’s
Alexander Volfovsky  alexvolf  1, 2a, 2b, 3c, 2c  functional analysis
Matthew Woolf  mwoolf  4a, 4b, 1, 3c, 2a, 3b, 3a, 2c, 2b manifolds, num analysis, diff equ.
Qian Sophia Zhang  qian  2a, 1, 2c, 2b, 3b

YSP: 7–8

Name (8)  @uchicago.edu  preferences  other topics
Emily Bargar*  ebargar  1, 2b, 3c, 4b  discrete math, algebra
Padraic James Bartlett*  padraic  2a, 3a, 4a, 3c (2c, 4b)  algebraic topology; topology
Sarah Constantin  sconsta1
John Dyer  dyer
Gabriel Gaster  eulerian  4b, 4a, 3c, 3b, 1  Galois thy, repn thy, number thy
Alice Mark  amark  1, 3b, 3c, 2a, 2b  algebra; number thy, algebraic geometry
James Marrone  jmar  4a, 2c/2a, 1  math. physics (chaos, info), Lie groups
Brian Taylor  bgtaylor  4, 1, 2, 3  algebraic number theory/algebra

* Do NOT pair these two with the same mentor! YSP: 9–10

Name (8) @uchicago.edu preferences other topics
Daniel Gardner  dbgardner
Drew A. Hudec  dhudec  4a, 3c, 3a, 4b  Algebraic geometry, homological algebra
Suyeon Khim  skhim  1, 2b, 3a, 4b, 3c, 3b
Grant Larsen  gel  4a, 1  algebraic number theory?
Jean Morrison  jeanm  4b, 1, 2b, 3b
Alexander J. Munk  amunk  2a, 1, 4b, 3b, 3c, 3a  probability (2a)
Prerna Nadathur  pnadathur  2c, 4a, 4b, 1, 3b  number thy, topology
Caitlin Waddle  kwaddle  1, 2b, 2a, 3a, 2c, 4b

YSP: 11–12

Name (6)  @uchicago.edu  preferences  other topics
Harry Altman  haltman  1, 3a, 2a, 3b, 4b, 4a, 2c, 2b
Corrin Clarkson  clarkson  2a, 3a, 4b (3c, 4a, 3b?)  algebra or topology
David Cohen  dcohen  1, 2a, 2b, 3a  analytic number thy, Apery’s thm
Lucas Culler  luccul
Jonathan Nieder jrnieder 2b, 1, 3b, 3c, 5, 2a number thy, math. phys, alg. geom
Sam Raskin  sraskin

COLLEGIATE SCHOLARS, OTHER FULL

Name (2)  @uchicago.edu preferences  other topics
Fan Fei Chong  ffclark  3a, 1, 4a  algebraic topology, comm. alg. dyn systems
Nicholas Ford  nford  1, 4, 3, 2 foundations (sets, cat. thy, model thy), algebra (gp thy)
Calvin Lin Zhiwei  linzw  3, 2 ,1